OpenStudy (kalijah13):

Ilya took one hour to drive from his apartment to Philips Arena and back. The return drive took 8 minutes less than the trip to the arena. If x represents the time it took Ilya to drive from his apartment to the arena, which equation would you use to determine how long he drove each way? A) x - 8 = 1 Eliminate B) 2x - 8 = 1 C) x - 8 = 60 D) 2x - 8 = 60

OpenStudy (kalijah13):

A SINGLE MEDALLION WILL BE AWARDED!

OpenStudy (kalijah13):

Please pardon the eliminate underneath A) for that was unintentional

OpenStudy (scarlettfarra2000):

An hour is 60 minutes. The trip to the arena + The return drive = 60 minutes. Since the return drive is 8 minutes shorter, if you added 8 minutes it would be the same length as the trip to the arena. The return drive + 8 minutes = The trip to the arena Let's represent these two equations with variables. x + r = 60 r + 8 = x What we do now is use substitution. Since r + 8 = x, we can replace x from the first equation with r + 8. x + r = 60 r + 8 + r = 60 2r + 8 = 60 2r = 52 r = 26 Then use r in an earlier equation to find a. r + 8 = x 26 + 8 = x so the answer should be D hope this helps you

OpenStudy (scarlettfarra2000):

@kalijah13